# Cost-sensitive computational adequacy of higher-order recursion in synthetic domain theory

### v2

@@ 4.4 Domains and admissibility
A subset of a domain $D$ is admissible when it contains $\bot_{D}$ and is closed under suprema of synthetic $\omega$-chains.
@@ 8.2 Adequacy
The proof of the fundamental lemma proceeds by induction on the derivation of terms. Details and the proof that formal approximation predicates $-\lhd_{\mathsf{U}X}e$ are admissible can be found in Appendix F; crucially we rely on the fact that $\mathsf{eval}(e,v)$ is a $\Sigma$-proposition.
@@ Appendix B Properties of predomains
Proposition B.8.
@@
Appendix C Properties of the denotational semantics
@@
Appendix D Properties of the computational semantics
@@
We may check that $P$ is admissible and proceed by fixed-point induction. Suppose that $\alpha\in P$ and that $\Phi_{\mathsf{eval}}(\alpha)(e,v){\downarrow}$ and $\Phi_{\mathsf{eval}}(\alpha)(e,v^{\prime}){\downarrow}$. We need to show that $v=v^{\prime}$. We proceed by cases on $\mathsf{out}(e)$.
@@
Appendix E Soundness of the denotational semantics
@@
Appendix F Proofs for the computational adequacy property
Proposition F.1.
@@ F.1 Admissibility
Proof F.4.
First, we have to show that $\bot\lhd_{\mathsf{U}\mathsf{F}A}e$. By definition, this means to show $f^{\sharp}(\bot)\sqsubseteq\mathsf{profile}(e;g)$ for all $f\mathrel{(A\Rightarrow\mathsf{adq})}g$. But this holds since $f^{\sharp}(\bot)=\bot$ and $\bot$ is the least element of $\mathsf{T}(A)$. Otherwise, let $d$ be a synthetic $\omega$-chain such that $d_{i}\lhd_{\mathsf{U}\mathsf{F}A}e$ for all $i:\omega$. We want to show that $\bigvee d\lhd_{\mathsf{U}\mathsf{F}A}e$, which is to show that $f^{\sharp}(\bigvee d)\sqsubseteq\mathsf{profile}(e;g)$ for all $f\mathrel{(A\Rightarrow\mathsf{adq})}g$. Since $f^{\sharp}(\bigvee d)=\bigvee(f^{\sharp}\circ d)$, this means to show $\bigvee(f^{\sharp}\circ d)\sqsubseteq\mathsf{profile}(e;g)$. By the universal property of the synthetic $\omega$-supremum, it suffices to show $f^{\sharp}(d_{i})\sqsubseteq\mathsf{profile}(e;g)$ for all $i:\omega$, but this is the assumption.
Proposition F.5.
@@
Proof F.7.
Because $\bot(a)=\bot$, we have that $\bot\lhd_{\mathsf{U}(A\to X)}e$. Suppose that $f_{i}\lhd_{\mathsf{U}(A\to X)}e$. We need to show that $\bigvee f\lhd_{\mathsf{U}(A\to X)}e$. Suppose that $a\lhd_{A}b$. We need to show that $(\bigvee f)~{}a\lhd_{\mathsf{U}X}e~{}b$. This follows from the characterization of synthetic $\omega$-suprema in function spaces (Proposition F.5) and the assumption that $-\lhd_{\mathsf{U}X}e$.
Proposition F.8.
@@
Proof F.9.
F.2 Fundamental lemma
@@
Proof F.11.
Let $f\mathrel{(\lhd_{A}\Rightarrow{\mathsf{adq}})}g$. We need to show that $(f^{\sharp}(\eta_{\mathsf{T}}(a)))\mathrel{\mathsf{adq}}(\mathsf{ret}(v);g)$. Computing the denotational semantics and applying Proposition F.1, it suffices to show that $(f~{}a)\mathrel{\mathsf{adq}}(g~{}v)$, which follows from our assumption.
Lemma F.12.
@@
    If $X=B\to Y$, suppose that $b\lhd_{B}v$. We need to show that $(f^{\sharp}(d))~{}b\lhd_{\mathsf{U}Y}(e;g)~{}v$. Unraveling the denotational semantics and the computational semantics (using Proposition D.6), it suffices to show $(\lambda d.~{}f~{}d~{}b)^{\sharp}~{}d\lhd_{\mathsf{U}Y}(e;\lambda d.~{}g~{}d~{}v)$, which follows from the inductive hypothesis and the assumption that $f\lhd_{\mathsf{U}(A\to(B\to Y))}g$.
Lemma F.14.
@@
Proof F.15.
Since $\mathsf{step}^{c}(e)\mapsto c,e$, the result holds by Proposition F.1.

### v3

@@ 4.4 Domains and admissibility
A subset of a domain $D$ is admissible when it is complete and contains $\bot_{D}$.
Remark 4.13.
It might be tempting to define admissible subsets as subsets of domains that are closed under the least element and synthetic $\omega$-joins, but it is unclear that this definition actually supports fixed-point induction in general. However this intuitive definition does suffice for all instances of fixed-point induction used in this paper.
Proposition 4.14.
A lower subset of a complete type is complete when it is closed under synthetic $\omega$-joins.
Corollary 4.15.
A lower subset of a domain is admissible when it is closed under the least element and synthetic $\omega$-joins.
@@ 8.2 Adequacy
The proof of the fundamental lemma proceeds by induction on the derivation of terms. Details and the proof that formal approximation predicates $-\lhd_{\mathsf{U}X}e$ are admissible can be found in the appendix Appendix G; crucially we rely on the fact that $\mathsf{eval}(e,v)$ is a $\Sigma$-proposition.
@@ Appendix B Properties of predomains
Appendix C Domains and admissibility
Proposition C.1.
The path relation need not coincide with the intrinsic order, but a path $\alpha:x\sqsubseteq^{\mathsf{p}}y$ always implies $x\sqsubseteq^{\circ}y$.
Proof C.2.
Suppose we have a path $l:\Sigma\to\Sigma$ whose boundary is determined by $x,y$. Fixing $f:\Sigma\to\Sigma$, we want to show that $f~{}x\to f~{}y$. But this follows by evaluating $\bot\sqsubseteq^{\circ}\top$ at the $\Sigma$-predicate $f\circ l:\Sigma\to\Sigma$.
Proposition C.3.
The invariant point $\infty$ is the top element of $\overline{\omega}$ with respect to the intrinsic preorder.
Proof C.4.
Observe that for any $i:\overline{\omega}$ we have a path $\alpha:i\sqsubseteq^{\mathsf{p}}\infty$ defined by $\alpha(\phi,n)=\phi\vee i(n)$. Thus we have isuperscriptsquare-$i\sqsubseteq^{\circ}\infty$ by Proposition C.1.
See 4.14
Proof C.5.
Suppose $S\subseteq A$ is downward closed and closed under synthetic $\omega$-joins. We must complete the following lifting problem:
$\omega$$S$$f$$\overline{\omega}$????
[…]
Thus it suffices to show that $\overline{\iota f}$ factors through $S$. Fixing $i:\overline{\omega}$, we want to show that $\overline{\iota f}(i)\in S$. By the assumptions on $S$, we have that $\bigvee(\iota f)=\overline{\iota f}(\infty)$ is in $S$ and that $a\in S$ whenever $a\sqsubseteq\overline{\iota f}(\infty)$. Thus it suffices to show $\overline{\iota f}(i)\sqsubseteq\overline{\iota f}(\infty)$, which follows from Proposition C.3.
Proposition C.6.
@@ Appendix C Domains and admissibility
Appendix D Properties of the denotational semantics
@@
Appendix E Properties of the computational semantics
@@
We may check that $P$ is admissible by means of Corollary 4.15 and proceed by fixed-point induction. Suppose that $\alpha\in P$ and that $\Phi_{\mathsf{eval}}(\alpha)(e,v){\downarrow}$ and $\Phi_{\mathsf{eval}}(\alpha)(e,v^{\prime}){\downarrow}$. We need to show that $v=v^{\prime}$. We proceed by cases on $\mathsf{out}(e)$.
@@
Appendix F Soundness of the denotational semantics
@@
Appendix G Proofs for the computational adequacy property
Proposition G.1.
@@ G.1 Admissibility
Proof G.4.
By Corollary 4.15 it suffices to show downward closure and closure under $\bot$ and synthetic $\omega$-joins.
[…]
    By definition this means to show $f^{\sharp}(\bot)\sqsubseteq\mathsf{profile}(e;g)$ for all $f\mathrel{(A\Rightarrow\mathsf{adq})}g$. But this holds since $f^{\sharp}(\bot)=\bot$ and $\bot$ is the least element of $\mathsf{T}(A)$.
[…]
    Let $d$ be a synthetic $\omega$-chain such that $d_{i}\lhd_{\mathsf{U}\mathsf{F}A}e$ for all $i:\omega$. We want to show that $\bigvee d\lhd_{\mathsf{U}\mathsf{F}A}e$, which is to show that $f^{\sharp}(\bigvee d)\sqsubseteq\mathsf{profile}(e;g)$ for all $f\mathrel{(A\Rightarrow\mathsf{adq})}g$. Since $f^{\sharp}(\bigvee d)=\bigvee(f^{\sharp}\circ d)$, this means to show $\bigvee(f^{\sharp}\circ d)\sqsubseteq\mathsf{profile}(e;g)$. By the universal property of the synthetic $\omega$-join, it suffices to show $f^{\sharp}(d_{i})\sqsubseteq\mathsf{profile}(e;g)$ for all $i:\omega$, but this is the assumption.
[…]
    Fixing $d^{\prime}\sqsubseteq d\lhd_{\mathsf{U}\mathsf{F}A}e$, we need to show that $d^{\prime}\lhd_{\mathsf{U}\mathsf{F}A}e$. Suppose that $f\mathrel{(A\Rightarrow\mathsf{adq})}g$. We need to show $f^{\sharp}(d^{\prime})\sqsubseteq\mathsf{profile}(e;g)$. By the characterization of the order on lifts Proposition 4.8, we suppose $f^{\sharp}(d^{\prime}){\downarrow}$ and show that $\mathsf{profile}(e;g){\downarrow}$ and that $f^{\sharp}(d^{\prime})=\mathsf{profile}(e;g)$. By assumption we know $d^{\prime}=\eta_{\mathbb{L}}(a^{\prime})$ and $f(a^{\prime})=c$ for some $a:\llbracket A\rrbracket$ and $c:\P\vee\mathbb{C}$. Since $d^{\prime}\sqsubseteq d$, we know $d^{\prime}=\eta_{\mathbb{L}}(a)$ for some $a$ such that $a^{\prime}\sqsubseteq a$. Consequently, we have $c=f(a^{\prime})\sqsubseteq f(a)$, but since $\P\vee\mathbb{C}$ is discrete (by Section 6 and the argument in Proposition 7.4), we have $f(a)=c=f(a^{\prime})$. Thus by the assumption that $d\lhd_{\mathsf{U}\mathsf{F}A}e$, we have $f^{\sharp}(d)=c\sqsubseteq\mathsf{profile}(e;g)$. Again by the discreteness of $\P\vee\mathbb{C}$ we have that $f^{\sharp}(d^{\prime})=f(a^{\prime})=c=\mathsf{profile}(e;g)$, as required.
Proposition G.5.
@@
Proof G.7.
Again we show downward closure and closure under $\bot$ and $\bigvee$.
[…]
    Because $\bot(a)=\bot$, we have that $\bot\lhd_{\mathsf{U}(A\to X)}e$ by the assumption that $\bot\lhd_{\mathsf{U}X}e$ for all $e$.
[…]
    Suppose that $f_{i}\lhd_{\mathsf{U}(A\to X)}e$. We need to show that $\bigvee f\lhd_{\mathsf{U}(A\to X)}e$. Suppose that $a\lhd_{A}b$. We need to show that $(\bigvee f)~{}a\lhd_{\mathsf{U}X}e~{}b$. This follows the fact that synthetic $\omega$-joins in function spaces are computed pointwise and the assumption that $-\lhd_{\mathsf{U}X}e~{}b$ is closed under $\bigvee$.
[…]
    Fix $f^{\prime}\sqsubseteq f\lhd_{\mathsf{U}(A\to X)}e$. To show that $f^{\prime}\lhd_{\mathsf{U}(A\to X)}e$, suppose that $a\lhd_{A}b$. We need to show that $f^{\prime}~{}a\lhd_{\mathsf{U}X}e~{}b$. By the premise, we have that $-\lhd_{\mathsf{U}X}e~{}b$ is a lower set, so it suffices to show $f^{\prime}~{}a\sqsubseteq f~{}a\lhd_{\mathsf{U}X}e~{}b$, which follow from the assumptions $f^{\prime}\sqsubseteq f$ and $f\lhd_{\mathsf{U}(A\to X)}e$.
Proposition G.8.
@@
Proof G.9.
G.2 Fundamental lemma
@@
Proof G.11.
Let $f\mathrel{(\lhd_{A}\Rightarrow{\mathsf{adq}})}g$. We need to show that $(f^{\sharp}(\eta_{\mathsf{T}}(a)))\mathrel{\mathsf{adq}}(\mathsf{ret}(v);g)$. Computing the denotational semantics and applying Proposition G.1, it suffices to show that $(f~{}a)\mathrel{\mathsf{adq}}(g~{}v)$, which follows from our assumption.
Lemma G.12.
@@
    If $X=B\to Y$, suppose that $b\lhd_{B}v$. We need to show that $(f^{\sharp}(d))~{}b\lhd_{\mathsf{U}Y}(e;g)~{}v$. Unraveling the denotational semantics and the computational semantics (using Proposition E.6), it suffices to show $(\lambda d.~{}f~{}d~{}b)^{\sharp}~{}d\lhd_{\mathsf{U}Y}(e;\lambda d.~{}g~{}d~{}v)$, which follows from the inductive hypothesis and the assumption that $f\lhd_{\mathsf{U}(A\to(B\to Y))}g$.
Lemma G.14.
@@
Proof G.15.
Since $\mathsf{step}^{c}(e)\mapsto c,e$, the result holds by Proposition G.1.